# Harnessing Environmental Memory with Reinforcement Learning in Open Quantum Systems

### v1

@@ I Introduction
Open quantum systems inevitably interact with surrounding environments, leading to decoherence, energy relaxation, and loss of distinguishability between quantum states [1, 2]. In many physical platforms from solid-state spins and superconducting qubits to nanophotonic structures and molecular systems the environment is structured or strongly coupled, and its correlations persist on timescales comparable to the intrinsic system dynamics. In such regimes, the reduced dynamics cannot be described by a simple Markovian semigroup and instead exhibits memory effects, revivals of coherence, and information backflow, i.e., genuinely non-Markovian behavior [3, 4, 5, 6, 7].
Over the past decade, non-Markovianity has been recognized as a useful resource rather than merely a nuisance. It can enhance quantum channel capacities and improve the performance of information-processing tasks [9, 10, 8], enable noise reduction beyond the reach of standard dynamical decoupling [11], and offer advantages in quantum metrology and sensing when suitably controlled [12, 13, 14]. Recent work has also highlighted the relevance of non-Markovianity for quantum machine learning, where environmental memory can boost the effective memory of quantum neural-network architectures [15]. These advances support a resource-theoretic view in which non-Markovian dynamics, viewed as multi-time quantum processes, can be harnessed and transformed by appropriate control operations [17, 16]
From a theoretical standpoint, controlling non-Markovian dynamics is challenging. Memory kernels, time-nonlocal master equations, or time-dependent decay rates such as $\gamma(t)$ in time-local formulations lead to control landscapes that are highly non-convex and history dependent [18, 19]. Standard optimal control theory (OCT) methods, which have been remarkably successful in Markovian and closed-system settings  [20, 21, 22], rely on the computation of gradients of a final-time cost functional with respect to control fields. In the presence of memory, these gradients become sensitive to numerical approximations and to the full past trajectory, which can lead to unstable optimization and trapping in local optima [23, 24]. This difficulty is amplified when the figure of merit is itself non-linear and temporally non-local, as is the case for most non-Markovianity measures.

### v2

@@ I Introduction
Open quantum systems inevitably interact with surrounding environments, leading to decoherence, energy relaxation, and loss of distinguishability between quantum states [1, 2]. In many physical platforms from solid-state spins and superconducting qubits to nanophotonic structures and molecular systems the environment is structured or strongly coupled, and its correlations persist on timescales comparable to the intrinsic system dynamics. In such regimes, the reduced dynamics cannot be described by a simple Markovian semigroup and instead exhibits memory effects, revivals of coherence, and information backflow, i.e., genuinely non-Markovian behavior [3, 4, 9, 5, 6, 7, 8].
Over the past decade, non-Markovianity has been recognized as a useful resource rather than merely a nuisance. It can enhance quantum channel capacities and improve the performance of information-processing tasks, for example by correlating memory effects with higher dense coding capacities in non-Markovian regimes  [10]. It has also been shown that temporal correlations from non-Markovian environments can be consumed to reduce noise beyond the reach of standard dynamical decoupling techniques [11]. Moreover, appropriately controlled non-Markovian effects can offer advantages in quantum metrology and sensing, enabling precision beyond the standard quantum limit [12, 13]. More generally, non-Markovian information backflow and system–environment memory can qualitatively alter open quantum dynamics with implications for quantum communication, simulation, and sensing protocols [12].
From a theoretical standpoint, controlling non-Markovian dynamics is challenging. Memory kernels, time-nonlocal master equations, or time-dependent decay rates such as $\gamma(t)$ in time-local formulations lead to control landscapes that are highly non-convex and history dependent [14]. Standard optimal control theory (OCT) methods, which have been remarkably successful in Markovian and closed-system settings  [15, 16], rely on the computation of gradients of a final-time cost functional with respect to control fields. In the presence of memory, these gradients become sensitive to numerical approximations and to the full past trajectory, which can lead to unstable optimization and trapping in local optima [17]. This difficulty is amplified when the figure of merit is itself non-linear and temporally non-local, as is the case for most non-Markovianity measures.
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the theoretical model of a driven two-level system coupled to a structured bosonic reservoir and recall the conditions under which non-Markovian memory effects arise. In Sec. III, we define the measure of non-Markovianity based on the Breuer–Laine–Piilo trace-distance criterion, which serves as the central figure of merit throughout this work. Section IV is devoted to a benchmark study using classical optimal control theory, where Powell and L-BFGS-B algorithms are employed to maximize the total non-Markovianity and to analyze the resulting limitations induced by memory-dependent control landscapes. In Sec. V, we introduce our reinforcement-learning framework, including the environment formulation, the reward design, and the implementation of the PPO and SAC algorithms for continuous quantum control. The numerical results and a detailed comparison between optimal control and reinforcement learning are presented and discussed in Sec. VI, highlighting the distinct strategies uncovered by each approach. Finally, Sec. VII summarizes our findings and outlines future directions for exploiting non-Markovian memory effects using data-driven quantum control methods.